OpenStudy (anonymous):

The rational function has a y-intercept of 7. What is the equation of this function?

OpenStudy (anonymous):

Okay so I'm in Algebra II and the book shows how to graph a rational function like this y=x^2+x-12/x^2-4, by finding the horizontal asymptote and the vertical asymptotes y=(x+4)(x-3)/(x+2)(x-2), then when the numerator equals 0, y=(0+4)(0-3)/(0+2)(0-2) = 3 so the y-intercept is (0,3), so my q is how do I use the graph to find the equation! Can someone please explain the steps the book doesn't show!!

OpenStudy (phi):

I guess work backwards. Your graph as a y intercept (0,7). you have a horizontal asymptote y=5 and a vertical asymptote x=-2

OpenStudy (anonymous):

Okay, I'm not sure where to go from here, I mean then i find x and y-intercepts, look for holes, but how would i put this in an equation??

OpenStudy (anonymous):

@phi do u know anyone who could teach this to me!?

OpenStudy (phi):

I remember that if you divide by 0 you get a vertical asymptote so to get an asymptote at x=-2 you want x+2 in the bottom , because -2+2 = 0 and we will divide by 0... creating an asymptote at x=-2

OpenStudy (anonymous):

Awesome now how to i get the numerator, same concept?? (x-5)/(x+2), that seems too simple..

OpenStudy (phi):

I am trying to remember what causes a horizontal asymptote I guess you don't remember ?

OpenStudy (anonymous):

So it says to find the horizontal asymptote of the graph of the rational function, compare the degree of the numerator m to the degree of the denominator n If m<n, the graph has a horizontal asymptote y=0 (the x-axis)v If m>n, the graph has no horizontal symptom If m=n the graph has horizontal asymptote y=a/b although i don't see how it relates...

OpenStudy (phi):

the degree is the highest power (or number of (ax+b) factors ) you have a horizontal asymptote (Agree ??) so we use If m=n the graph has horizontal asymptote y=a/b

OpenStudy (anonymous):

m=n y=a/b just to add to the last one a is the coefficient of the term of greatest degree in the numerator and b is the coefficient of the term of the greatest degree in the denominator

OpenStudy (phi):

we know we have (x+2) in the bottom the "b" they are talking about is the number in front of the x i.e. b=1 up top we want (a x + c ) then we will get a horizontal asymptote at y= a/b = a/1 or just y=a what should a be ?

OpenStudy (anonymous):

5?

OpenStudy (phi):

so now we have (5x + c) / (x+2) we know that will give us a vertical asymptote at x= - 2 (divide by 0) and according to the rules you just posted, it will give a horizontal asymptote at y = 5/1 = 5 both of those match your graph. we still need to figure out what c is ? I would use the info on the y intercept.

OpenStudy (anonymous):

Okay so 7 is the y-intercept

OpenStudy (anonymous):

Is that it 5x+7/ x+2 ? @phi

OpenStudy (phi):

close. but so far we have y = (5x+c)/(x+2) we know when x is 0, y = 7 (definition of y-intercept) so replace y with 7 and x with 0 in the equation, and solve for c

OpenStudy (anonymous):

7=5(0)+c/0+2 7=c/2 c= 14??

OpenStudy (phi):

yes

OpenStudy (anonymous):

Okay so the equation is y=5x+14/x+2

OpenStudy (phi):

yes, but I would use parens y = (5x+14)/(x+2) here is a graph (using geogebra ) of that equation and it looks like a good match

OpenStudy (anonymous):

It does look correct!! although the top 5x+14, do u try to factor that out or just leave as is, cause i think its fine :)

OpenStudy (phi):

you can't simplify 5x+14 if it were (6x+14) you could write it as 2(3x+7) for example but 5 and 14 have no common factors. the answer is just \[ y = \frac{5x+14}{x+2} \]

OpenStudy (anonymous):

Right, oops forgot! now i understand, THANK YOU so much i might be the only kid in the class who can do the reverse AWESOME!!! can't tell u how much i learned :) its great u don't just tell concept but show how to apply it, never would have gotten it without u :D